In a particular small town, the population in 1995 was 1200. Ever since then the population has grown by 4% every year. If conditions stay the same, what is the approximate expected population in the year 2000?
Answer by stanbon(57387)

In a particular small town, the population in 1995 was 1200. Ever since then the population has grown by 4% every year. If conditions stay the same, what is the approximate expected population in the year 2000?
------------------------------------------------------------
Population in 200= [1200+0.04*1200]^5 = 1200*1.04^5 = 1200*1.2167
----
= 1459.99
---
I guess you could round up: 1460
=======================================================
Cheers,
Stan H.
===============